### v2

# MetMamba: Regional Weather Forecasting with Spatial-Temporal Mamba Model

Abstract
Deep Learning based Weather Prediction (DLWP) models have been improving rapidly over the last few years, surpassing state of the art numerical weather forecasts by significant margins. While much of the optimization effort is focused on training curriculum to extend forecast range in the global context, two aspects remains less explored: limited area modeling and better backbones for weather forecasting. We show in this paper that MetMamba, a DLWP model built on a state-of-the-art state-space model, Mamba, offers notable performance gains and unique advantages over other popular backbones using traditional attention mechanisms and neural operators. We also demonstrate the feasibility of deep learning based limited area modeling via coupled training with a global host model.
1 Introduction
Weather forecasting has attracted increasing attention as climate change brought more heat waves, tropical cyclones, heavy rainfalls and other extreme weather events, affecting millions of people worldwide [1] [2]. Researchers have devised sophisticated numerical schemes and equations to capture complex weather dynamics to improve forecast accuracy[3]. However, its immense computational complexity often necessitates the use of large scale compute clusters, incurring notable energy cost and long processing time, making ensemble forecasts, critical for predicting such events[4], expensive or infeasible.
Trained on European Centre for Medium-range Weather Forecast (ECMWF)’s reanalysis product ECMWF Reanalysis v5 (ERA5) dataset, deep learning based weather prediction (DLWP) models have shown promising performance, with FourCastNet[5] being the first data-driven model that directly competes with ECMWF’s Integrate Forecasting System (IFS), followed by a series of other models[6][7][8][9] that outperforms IFS in various metrics [10], these models demonstrates excellent ability in capturing global weather trends, with a fraction of the computation requirement.
Though a coarser grid global forecast is crucial for gaining meaningful insights into longer horizon outlooks, extreme weather events often come in the form of convection, which are sub-grid dynamics not reflected in coarser grid data, even with advancements in fine grid global forecast, such events also requires localized tunning[11], making limited area model (LAM) pivotal in many applications. We conduct this research within the LAM context, partly due to its greatly reduced compute, but also aim to explore the feasibility of deep learning for Limited Area Modeling (LAM).
The Mamba model[12] exhibits remarkable performance with linear complexity and unique advantages on contiguous data, its vision variants[13] also features an effective receptive field similar to that of convolution networks, making it a promising alternative to the transformer dominated DLWP models, enabling deeper models and finer resolutions to be used, capturing more complex dynamics.
2 Related Work
2.1 Neural Operators
Neural Operators have shown great efficiency in solving partial differential equations, effective in a range of synthetic and physics-based tasks[14][15]. It has also shown similar efficiency in vision tasks as a token mixer when used in a Vision Transformer (ViT) like architecture[16]. FourCastNet builds on this model and delivers the first DLWP model capable of matching the performance of leading numerical weather prediction (NWP) system, ECMWF’s IFS. Further developments have been made by adopting spherical harmonics to better address the characteristics of global weather[17].
2.2 Transformers
Transformers, known for its ability to associate information in various parts of the input data via the attention mechanism[18], are initially designed for natural language processing tasks, have quickly gain wide adoption in computer vision with Vision Transformers[19]. Its quadratic complexity (1) however, poses as a significant hurdle in its application, particularly in vision. Different approaches to alleviate this issue have been proposed, most popular among them is Swin-Transformer[20], which uses a switching window mechanism to reduce the number of required tokens to achieve a global receptive field. Sub-quadratic approximations of attention like Linformer[21] are also a viable surrogate. In the DLWP space, where the compute efficiency is of a lesser concern, the use of transformers is ubiquitous[7][8][22][9][23], favoured by its modeling power. But as noted by [23], the patching step which are used to reduce the number of tokens also leads to degradation in performance.
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
2.3 Mamba Model
Mamba[12] model is a class of State Space Models (SSMs), which is originally designed to process continuous signals, bearing close similarity in its formulation to RNNs. (Eq.(2)) Its primary contribution, aside from parallelization and hardware level optimization, is enabling SSMs to process inputs in a time dependent manner, making it a viable alternative to attention [24] Modifying it for vision use cases involves transferring its 1D-Scan pattern to a 2d space[13][25], which achieves better performance than other leading vision transformer models.
$h_{t}$ $=\overline{A}h_{t-1}+\overline{B}x_{t}$ (2)
$y_{t}$ $=Cx_{t}$ (3)
2.4 Limited Area Modeling
The paradigm of using Lateral Boundary Conditions (LBCs) from a global model in Limited Area Modeling (LAM) is a common practise in running more location specific limited area models.
The usage of LBCs, specifically from forecast fields of a global model, has yet to be tried in the DLWP context, with neural-LAM[26] integrating ground truth field as LBCs. The main objectives of our paper are to evaluate a high performance LAM with LBCs from a global model, optimizing the coupling of the two models, and quantifying performance gains in a more realistic scenario.
3 Methodology
3.1 Dataset
The dataset we used in this paper is the ECMWF Reanalysis v5 (ERA5)[27][28] data. We collected ERA5 data from the year 2006 through 2022 in 6 hour intervals, using the area spanning: 11.25°11.25°11.25  °~60°60°60  °N, 73°73°73  °~135.75°135.75°135.75  °E. We use year 2006 to 2020 for training, 2021 for validation and 2022 for testing.
For the lateral boundary condition, we chose the latest iteration of FourCastNet (SFNO), distributed with Earth2Studio[29]. Among models with wide adoption and benchmarks, it is the only one with a continuous 6 hour rollout scheme, specific humidity as its vapor related variable and sufficient years outside of the training set for LBC generation. The LBCs are produced by initializing the model with ERA5 global data from 2018 onward, making a relatively smaller (2018~2021) for LBCs adaptation training, avoiding the years where the global model was trained on. See Appendix A for details regarding the dataset used.
Figure 1: (a) general model architecture: the DLWP-LAM takes primary inputs: two initial conditions $IC_{0}$ at initialization time $t_{0}$ and $IC_{-1}$ at previous (6 hours before) initialization time $t_{0-6h}$, and auxiliary inputs: constants from Digital Elevation Model (DEM) and elapsed time, they are processed by a DEM specific embedder, a conditioning module or by simple concatenation. The input tensor is then processed by $N$ blocks. The output is decoded with a simple linear layer and pixel shuffle operation. (b) DEM Embedder: an embedder that processes different information from the DEM accordingly. (c) AdaLN Block: a block that integrates elapsed year time via the use of adaLN, this will regulate all blocks in the process step.
3.2 Models
To illustrate the performance advantages of a Mamba based backbone for weather, we designed a general network architecture with three different types of block as their core processing unit: Swin, Swin+AFNO and Mamba. The models share similar parameter count and depth. The models all take two initial conditions and relevant boundary conditions as input. For model implementation details, ablation studies of relevant modules, see Appendix B
General Architecture
We applied the same design to all models, using a 3D convolution module encoding the two input initial conditions, with Mamba-3D being the exception as it can natively process spatial temporal data. Each model features the same Digital Elevation Model(DEM) embedder and AdaLN conditioning to process auxiliary topographical and temporal information. No down or up sampling are used, with model embedding size and latent space resolution staying the same throughout the main processing blocks, which we have found to be beneficial to the general modeling capability. See Figure 1
Swin Block
This is a Swin-Transformer-V2 block used in Swin-Transformer-V2[30] and also by FuXi[8], with no modifications, we refer to their paper for implementation details. LAM using this block is referred to as MetSwin.
Swin-AFNO Fusion Block
Adaptive Fourier Neural Operator, despite its efficiency, has rather limited skill ceiling in our earlier LAM based experiments, consistently performing worse than all other types of model. This is also reflected in benchmarks by several global DLWP papers. It is also prune to generate defects and even model collapse[17].
We design a dual branch model that mixes latent information between Fourier Neural Operator and SwinV2 branches using a FFT cross attention mechanism [31], achieving better results than using either of the single branch models, while alleviating the limitation of AFNO in the limited area setting. LAM using this block is referred to as MetSwinAFNOFusion.
Mamba-3D Block
We adopt the cross-scan and cross-merge methods used by VMamba[13], and extend these methods to the 3 dimensional space, to better integrate information from spatial and temporal dimensions, similarly, a 3D depth-wise convolution module is also used. We delineate the design choices of Mamba-3D in Appendix B.2.3 LAM using this block is referred to as MetMamba or MetMamba-3D, its 2 dimensional counterpart with basic 2D spatial scan pattern is referred to as MetMamba-2D.
DEM encoder
We encode different static information including DEM data used in the Weather Research and Forecast (WRF) Model, latitude and longitude coordinates into the model using a specific DEM encoder. We separate these data based on their datatype, categorical data (e.g. soil class, land sea mask) is handled by specific embedding layers, and numerical data is processed by simple linear layers.
AdaLN conditioning
Following FiLM[32] and Stormer [23], we encode 1 dimensional time data: elapsed year time to better regulate model against seasonal variations in weather dynamics.
Figure 2: Lateral Boundary Condition merging scheme, used in LBC adaptation training and inference rollout
3.3 Lateral Boundary Condition
[…]
3.4 Loss Function
We use Mean Squared Error for the loss function. The loss is calculated by summing and averaging all loss values for each weather variable across all pressure levels, auto-regressive rollout time steps on grid points within the area constrained by the Lateral Boundary Condition.
It is weighted generally following the scheme of GraphCast, with height-based weights focusing on surface and near ground variables, and latitude-based weights assigning higher multipliers to grid points with lower latitude. All weights are fixed in all experiments presented in this paper.
$\mathcal{L}=\frac{1}{N_{\text{Rollout}}}\sum_{t=t_{\text{Init}}+1}^{t_{\text{Init}}+N_{\text{Rollout}}}\frac{1}{\left|\mathbb{G}\right|}\sum_{v\in\mathbb{G}}\sum_{i=1}^{S}\omega_{i}\left(\hat{X}^{t}_{v,i}-X^{t}_{v,i}\right)^{2}$ (4)
where
  • $t_{\text{Init}}$ is the initial time step the forecast starts from,
  • $N_{\text{Rollout}}$ is the number of rollout steps used during training,
  • $\mathbb{G}=\{1,\dots,N\}\setminus\mathbb{B}$ is the set of grid nodes not on the boundary,
  • $\omega_{i}$ is a weight associated with the variable $i$. This varies depending on variable category, pressure level and latitude.
3.5 Training Curriculum
Pre-Training
The model is trained first to predict the weather state of the next time step for 30 epochs each consisting of 10957 gradient updates with the AdamW optimizer, which follows a linear warmup and cosine annealing schedule.
LBC Adaptive Auto-Regressive Training
As noted by [33], the convergence of the LBCs with the LAMs is the more important factor when producing high fidelity results, than the well-posedness of the LBCs. We thus introduced these LBCs into the training process. LBCs sampled from the global model FourCastNet(SFNO) are merged with the predicted results of the LAM and used in turn as input to the LAM. This process replaces the auto-regressive training step usually found in the global DLWPs. We set this auto-regressive step count to be 4.
3.6 Evaluation
We evaluate all LAM models, FourCastNet(SFNO) and relevant ablation studies with initial conditions from test year 2021. In the total 1439 initial conditions, only 1418 of them are used, with the first 1 and last 20 initial conditions reserved for model initialization and forecast truncation. The models are run and evaluated for 20 steps. Only Root Mean Squared Error (RMSE) is provided as a performance metric for simplicity. We note that this metric is calculated following the convention by WeatherBench[34], with averaging as the last step of the calculation, latitude weighting scheme is not used.
3.7 Environment
The model and revelant training, evaluation and inference apparatus are developed using PyTorch [35], xarray [36].
These models are trained on two Nvidia A100-40GB cards utilizing PyTorch’s Fully Sharded Distributed Parallel, gradient checkpointing and Automatic Mixed Precision in bf16 datatype.
4 Results
Figure 3: Root Mean Square Error (RMSE) of headline variables, comparing against ground truth ERA5 in 1418 5-day forecasts, evaluated on the test set of year 2022, for 3 types of DLWP-LAMs discussed in this paper. We omit variable $10u$ here for its similarity with the variable $10v$.
[…]
The Mamba based model has a notable performance margin over the other two LAMs, able to match or surpass on all 69 of the predicted variables from its global host model, while simultaneously resolving more detail than its counterparts.
We also note that the DLWP-LAMs are prune to generating artifacts at long lead times, which is a known phenomenon in both constrained DLWP-LAMs and unconstrained global DLWP models, a manifestation of the inherent flaws in the model architecture, i.e. checkerboard artifacts in Swin-Transformer based models and complete collapse of AFNO based models. This is alleviated by more robust auto-regressive training. We note that the Mamba based model have no noticeable artifacts, hinting at its unique suitability in weather forecasting.
For detailed evaluations, forecast examples and spectra analysis, see Appendix C.
5 Discussion
Mamba for Weather
The Mamba block, when used as a drop-in replacement in our existing LAMs offers a notable performance uplift. This can be attributed to its ability to both generate global receptive field and associate nearby tokens more efficiently than convolution networks. A reduced direct association between disjointed and distant tokens than vision transformers, which we suspect is less common in the field of meteorology, could also be a contributing factor. Its mathematical formulation being more suited to process continuous signals is another valid hypothesis. We also note that the cross scan pattern for VMamba and MetMamba is natively transferable and more suitable to a global geometry.
Figure 4: 4 steps of auto-regressive training: Normalized Root Mean Square Error (Norm. RMSE) of headline variables, comparing against FourCastNet (SFNO) in 1418 5-day forecasts, evaluated on the test set of year 2022. We omit variable $10u$ here for its similarity with the variable $10v$. The LAM model shows improvements on long lead times as the LBC-Merging auto-regressive training progresses.
Training for a Limited Area Model
Auto-regressive training is a universally adopted practise in almost all of the global DLWP models. Mimicking the model coupling step in numerical LAM systems, we train the DLWP-LAMs with LBC data merged from the host global model, exposing them to the characteristics of the LBCs, significantly improving the compatibility between the two, enhancing the forecast capability of the LAM. The effects of this adaptive training process is shown in Figure 4.
Auxiliary Information Encoding
: Previous models have used simple linear layers to process elevation, soil, land-sea mask and other time related information. We use specialized modules according to process each type of data, See Appendix B.2 for effects of these modules.
Performance Headroom
[…]
Acknowledgments
The ECMWF Reanalysis v5 (ERA5) [27] [28] dataset is used in this paper to train various models, it is retrieved via Copernicus Climate Data Store.
We thank European Centre for Medium-Range Weather Forecasts (ECMWF) and its C3S service for their efforts in collecting, archiving, and disseminating the data, without which studies like this would not have been possible.
Appendix A Data Details
Type Variable Name Description ECMWF Parameter ID Role
Atmospheric z@--- Geopotential (at pressure level ---) 129 Input/Ouptut
Atmospheric t@--- Temperature (at pressure level ---) 130 Input/Ouptut
Atmospheric u@--- u component wind (at pressure level ---) 131 Input/Ouptut
Atmospheric v@--- v component wind (at pressure level ---) 132 Input/Ouptut
Atmospheric q@--- Specific humidity (at pressure level ---) 133 Input/Ouptut
Surface msl Mean sea level pressure 151 Input/Ouptut
Surface 10u 10 metre u wind component 165 Input/Ouptut
Surface 10v 10 metre v wind component 166 Input/Ouptut
Surface 2t 2 metre temperature 167 Input/Ouptut
Static Longitude Sine normalized n/a Input
Static Longitude Cosine normalized n/a Input
Static Latitude Sine normalized n/a Input
Static LANDMASK Land sea mask n/a Input
Static LU_INDEX Land use index n/a Input
Static SNOALB Maximum snow albedo (MODIS) n/a Input
Static SANDFRAC Sand content in soil n/a Input
Static HGT_M Topography height (GMTED2010) n/a Input
Static SCB_DOM Dominant soil class n/a Input
Static SCT_DOM Dominant surface cover type n/a Input
Clock Local Time of Day Sine normalized n/a Input
Clock Local Time of Day Cosine normalized n/a Input
Clock Elapsed Year Progress Sine normalized (1D) n/a Input
Clock Elapsed Year Progress Cosine normalized (1D) n/a Input
Clock Forecast lead time Fixed value of 6 (z-score normalized) n/a Input
Table 1: Pressure level variables, surface variables and auxiliary used and predicted by the model and their notation style,
In this appendix we give more details on the dataset and other auxiliary features used in our experiments.
Data collection, notation and processing
We collected ERA5[27][28] data with ECMWF’s Copernicus Climate Change Service (cdsapi), targeting the area spanning: 11.25°11.25°11.25  °~60°60°60  °N, 73°73°73  °~135.75°135.75°135.75  °E, atmospheric variables: geopotential z@---, temperature t@---, $u$-component wind u@---, $v$-component wind v@--- specific humidity q@---, 10 metre $u$ wind component 10u, 10 metre $v$ wind component 10v, mean sea level pressure msl, 2 metre temperature 2t. We calculate the mean and standard deviation for all variables throughout the training subset, and perform z-score normalization for all of the variables prior to model input. For all variables and notations, see Table 1.
Topographical Data
Digital Elevation Model data is sampled from the Weather Research and Forecast (WRF) model, containing land-sea mask, soil type, snow albedo etc. Positional information is added by calculating the sine and cosine normalized latitude and longitude. (See Eq.(A) and Eq.(6)) Time of day is added as a group of 2 dimensional sine and cosine normalized local time. Time of year is added as a group of 1 dimensional sine and cosine normalized current UTC time (in seconds) divided by total seconds per year.
Lateral Boundary Conditions
The choice of Lateral Boundary Conditions, is subject several constraints. To be eligible for LBC adaptive training, the global DLWP model has to: 1) share the same variables with the LAM; 2) leave sufficient years out of the training set, such that a realistic LBCs can be reproduced; 3) use a consistent autoregressive rollout strategy; As such we have chosen FourCastNet(SFNO) in this paper as it satisfies all criteria.
$C=\frac{\cos\left(C\times\frac{\pi}{180}\right)-\frac{1}{N}\sum_{i=1}^{N}\cos\left(C_{i}\times\frac{\pi}{180}\right)}{\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\cos\left(C_{i}\times\frac{\pi}{180}\right)-\frac{1}{N}\sum_{j=1}^{N}\cos\left(C_{j}\times\frac{\pi}{180}\right)\right)^{2}}}$ (5)
$C=\frac{\sin\left(C\times\frac{\pi}{180}\right)-\frac{1}{N}\sum_{i=1}^{N}\sin\left(C_{i}\times\frac{\pi}{180}\right)}{\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\sin\left(C_{i}\times\frac{\pi}{180}\right)-\frac{1}{N}\sum_{j=1}^{N}\sin\left(C_{j}\times\frac{\pi}{180}\right)\right)^{2}}}$ (6)
Appendix B Model Details
We delineate the design and implementation details of the our MetMamba model in this section. The main difference to two other types of DLWP-LAMs is MetMamba’s ability to process spatial-temporal inputs natively, without the need for an additional 3D convolution encoder. Aside from sharing the same structure, auxiliary information encoder and decoder, all DLWP-LAMs share the same embedding size and similar parameter count. The SwinV2 based LAM serves as a baseline, the AFNO-Swin based LAM serves as a incremental update with relatively higher efficiency and performance. See Table 2.
For the detailed design of the SwinV2 block, we refer to the Swin-Transformer-V2 paper [30]. For AFNO-Swin Fusion block, we follow the dual branch design from Cross-ViT [31], with one AFNO block for each of the four model stages, the branches are mixed together with FFT truncation operations.
B.1 Model Design
Figure 5: Met-Mamba Architecture (a) Met-Mamba model architecture, the model takes primary inputs: two initial conditions $IC_{0}$ at initialization time $t_{0}$ and $IC_{-1}$ at previous (6 hours before) initialization time $t_{0}-6h$, and auxiliary inputs: DEM masks, elapsed time etc, processed by a specific embedder, linear layer, conditioning module or by simple concatenation. The spatial-temporal tensor is than processed by $N$ Mamba3D blocks, conditioned by elapsed time (year). The output is decoded with a simple linear layer and pixel shuffle operation. (b) AdaLN-Mamba3D Block, a block that utilizes depth-wise 3D convolution and mamba’s selective scan to achieve token (spatial, temporal) mixing and channel mixing. The block integrates seasonal variation via the use of adaLN. (c) SS3D operator, a module that flattens and rearranges input spatial-temporal tensor to achieve different memory layouts (scan routes) for Mamba’s selective scan to associate information on different directions.
model embed size patch size depth # of params samples/s
MetSwin 192 (1, 1) 22 19.4M 6.02
MetAFNOSwinFusion 192 (1, 1) 16+4 18.1M 5.93
MetMamba-2D 192 (1, 1) 22 21.8M 3.10
MetMamba-3D 192 (1, 1) 22 20.0M 1.72
Table 2: DLWP-LAM Models used in this paper, note that the parameter regression of MetMamba-3D is due to its reduced embedding size of $(H,W)$, rather than the $(C,H,W)$ used for other models. Static and auxiliary data are used only as input (static and time variant)
Scan routes
To flatten a spatial-temporal block of $(T,H,W)$ for Mamba-3D to scan, there exists 6 permutations of this block and 6 corresponding traversal routes:
  1. (1)
    Horizontal-Vertical-Temporal ($(T,H,W)$) : $K$ row first 2D scans;
  2. (2)
    Vertical-Horizontal-Temporal ($(T,W,H)$) : $K$ column first 2D scans;
  3. (3)
    Horizontal-Temporal-Vertical ($(H,T,W)$) : Row first cross temporal scan;
  4. (4)
    Vertical-Temporal-Horizontal ($(W,T,H)$) : Column first cross temporal scan;
  5. (5)
    Temporal-Horizontal-Vertical ($(H,W,T)$) : Temporal first psuedo 2D row scan;
  6. (6)
    Temporal-Vertical-Horizontal ($(W,H,T)$) : Temporal first psuedo 2D column scan;
where $K$ is the number of initial conditions, each route has a respective reverse route, making a total of 12 routes. These routes are designed to enable the model to build a effective receptive field (ERF) that can associate information with nearby patches across both spatial and temporal dimensions. We have observed large performance discrepancies when using different groupings of these routes, with the best combination being routes Horizontal-Vertical-Temporal and Temporal-Vertical-Horizontal. Note that while this grouping is the best route grouping based on limited experiments, heuristics and speculative convergence behaviour, a better grouping may still exist.
Usage of SS3D
There are several ways we can apply this SS3D block, we had limited experiments with the following 3:
  1. (1)
    Use as a drop-in encoder for encoding two input initial conditions;
  2. (2)
    Use in backbone, to process spatial-temporal relations for inputs with two initial conditions;
  3. (3)
    Use in backbone, to process spatial relations for inputs with one initial condition;
And we have found that, the best performance is yielded by model using a full spatial-temporal 3D backbone, we suspect this scan pattern and implementation is insufficient to exploit the spatial relationship but good enough to spatial temporal information.
B.2 Ablation Study
We replaced modules in the encoding and backbone stages of the model to observe their effect on forecasting capabilities.
B.2.1 DEM embedder
We replace the DEM embedder with a all-linear module, slight degradation of performance on 2t and variables higher than 150 hPa is observed, see Figure 6
Figure 6: Root Mean Squared Error (RMSE), MetMamba-2D when comparing against the same model but with the DEM embedder replaced with regular linear layers
B.2.2 AdaLN Conditioning
We removed the adaLN module from the model and supply the elapsed year time related information via expanded 2 dimensional tensors. Performance regression can be observed over all variables, see Figure 7
Figure 7: Root Mean Squared Error (RMSE), MetMamba-2D when comparing against the same model but with the elapsed year time information supplied as a expanded channel, rather than as through adaLN
B.2.3 Mamba3D
We replace spatial-temporal Mamba3D blocks with vanilla Mamba2D, which comes natively with VMamba, and observed a slight performance degradation across all variables, with geopotential z and variables above 150 hPa seeing the most regression, see Figure 8
Figure 8: Root Mean Squared Error (RMSE), MetMamba (2D) when comparing against the MetMamba (3D) with the same overarching architecture
Appendix C Detailed Results
In this appendix showcase additional results from the MetMamba model, when compared against FourCastNet(SFNO) and ERA5.
C.1 RMSE
The DLWP-LAMs are able to surpass or match the performance of host FourCastNet (SFNO) with significantly less training data, with MetMamba being the only one beats SFNO with notable margins on almost all variables with a few exceptions of v@50 and v@100, see Figure 9
Figure 9: Root Mean Squared Error (RMSE), 3 types of LAM comparing against their global host model FourCastNet (SFNO)
C.2 Example Forecasts
Here we showcase forecast results at lead time 24, 72 and 120 hours, initialized with ERA5 data at 2022-6-30 00:00:00 UTC. See Figure 1017
Figure 10: Example Forecasts: 2t, 3 types of LAM comparing against their global host model FourCastNet (SFNO) initialized at 2022-6-30 00:00:00 UTC, showing forecast at 24, 72 and 120 hour lead times
Figure 11: Example Forecasts: 10u, 3 types of LAM comparing against their global host model FourCastNet (SFNO) initialized at 2022-6-30 00:00:00 UTC, showing forecast at 24, 72 and 120 hour lead times
Figure 12: Example Forecasts: 10v, 3 types of LAM comparing against their global host model FourCastNet (SFNO) initialized at 2022-6-30 00:00:00 UTC, showing forecast at 24, 72 and 120 hour lead times
Figure 13: Example Forecasts: msl, 3 types of LAM comparing against their global host model FourCastNet (SFNO) initialized at 2022-6-30 00:00:00 UTC, showing forecast at 24, 72 and 120 hour lead times
Figure 14: Example Forecasts: t@850, 3 types of LAM comparing against their global host model FourCastNet (SFNO) initialized at 2022-6-30 00:00:00 UTC, showing forecast at 24, 72 and 120 hour lead times
Figure 15: Example Forecasts: q@700, 3 types of LAM comparing against their global host model FourCastNet (SFNO) initialized at 2022-6-30 00:00:00 UTC, showing forecast at 24, 72 and 120 hour lead times
Figure 16: Example Forecasts: v@850, 3 types of LAM comparing against their global host model FourCastNet (SFNO) initialized at 2022-6-30 00:00:00 UTC, showing forecast at 24, 72 and 120 hour lead times
Figure 17: Example Forecasts: u@850, 3 types of LAM comparing against their global host model FourCastNet (SFNO) initialized at 2022-6-30 00:00:00 UTC, showing forecast at 24, 72 and 120 hour lead times
C.3 Spatial Distribution of Error
Here we showcase the (averaged on test year 2022) spatial distribution of error (RMSE) of MetMamba versus FourCastNet(SFNO) across all headline variables. Both DLWP models are evaluated against ERA5.
For side by side comparisons, see Figure 1826.
For comparisons of error differential, see Figure 2735
In the illustrated examples, we observed degraded performance along the edges and in long lead times, suggesting the LBCs still played a role in limiting the performance of LAM. We also note the performance improvements on mountainous areas in a majority of variables, with the exception of u@850 and v@850. Large error differential spikes can also be found in north-western China, where the climate tends to be drier, this could be a manifestation of the overestimation of water contents in the Mamba based model.
[…]
Figure 18: Root Mean Squared Error (RMSE) of FourCastNet (SFNO) and MetMamba of variable 10u, when both compared against ERA5. Darker red indicates worse performance observed at this grid point.
[…]
Figure 19: Root Mean Squared Error (RMSE) of FourCastNet (SFNO) and MetMamba of variable 10v, when both compared against ERA5. Darker red indicates worse performance observed at this grid point.
[…]
Figure 20: Root Mean Squared Error (RMSE) of FourCastNet (SFNO) and MetMamba of variable 2t, when both compared against ERA5. Darker red indicates worse performance observed at this grid point.
[…]
Figure 21: Root Mean Squared Error (RMSE) of FourCastNet (SFNO) and MetMamba of variable msl, when both compared against ERA5. Darker red indicates worse performance observed at this grid point.
[…]
Figure 22: Root Mean Squared Error (RMSE) of FourCastNet (SFNO) and MetMamba of variable z@500, when both compared against ERA5. Darker red indicates worse performance observed at this grid point.
[…]
Figure 23: Root Mean Squared Error (RMSE) of FourCastNet (SFNO) and MetMamba of variable q@700, when both compared against ERA5. Darker red indicates worse performance observed at this grid point.
[…]
Figure 24: Root Mean Squared Error (RMSE) of FourCastNet (SFNO) and MetMamba of variable t@850, when both compared against ERA5. Darker red indicates worse performance observed at this grid point.
[…]
Figure 25: Root Mean Squared Error (RMSE) of FourCastNet (SFNO) and MetMamba of variable u@850, when both compared against ERA5. Darker red indicates worse performance observed at this grid point.
[…]
Figure 26: Root Mean Squared Error (RMSE) of FourCastNet (SFNO) and MetMamba of variable v@850, when both compared against ERA5. Darker red indicates worse performance observed at this grid point.
[…]
Figure 27: Root Mean Squared Error (RMSE) difference between FourCastNet (SFNO) and MetMamba of variable 10u, when both compared against ERA5. Blue indicates better performance observed at this grid point, red indicates worse performance observed at this grid point.
[…]
Figure 28: Root Mean Squared Error (RMSE) difference between FourCastNet (SFNO) and MetMamba of variable 10v, when both compared against ERA5. Blue indicates better performance observed at this grid point, red indicates worse performance observed at this grid point.
[…]
Figure 29: Root Mean Squared Error (RMSE) difference between FourCastNet (SFNO) and MetMamba of variable 2t, when both compared against ERA5. Blue indicates better performance observed at this grid point, red indicates worse performance observed at this grid point.
[…]
Figure 30: Root Mean Squared Error (RMSE) difference between FourCastNet (SFNO) and MetMamba of variable msl, when both compared against ERA5. Blue indicates better performance observed at this grid point, red indicates worse performance observed at this grid point.
[…]
Figure 31: Root Mean Squared Error (RMSE) difference between FourCastNet (SFNO) and MetMamba of variable z@500, when both compared against ERA5. Blue indicates better performance observed at this grid point, red indicates worse performance observed at this grid point.
[…]
Figure 32: Root Mean Squared Error (RMSE) difference between FourCastNet (SFNO) and MetMamba of variable q@700, when both compared against ERA5. Blue indicates better performance observed at this grid point, red indicates worse performance observed at this grid point.
[…]
Figure 33: Root Mean Squared Error (RMSE) difference between FourCastNet (SFNO) and MetMamba of variable t@850, when both compared against ERA5. Blue indicates better performance observed at this grid point, red indicates worse performance observed at this grid point.
[…]
Figure 34: Root Mean Squared Error (RMSE) difference between FourCastNet (SFNO) and MetMamba of variable u@850, when both compared against ERA5. Blue indicates better performance observed at this grid point, red indicates worse performance observed at this grid point.
[…]
Figure 35: Root Mean Squared Error (RMSE) difference between FourCastNet (SFNO) and MetMamba of variable v@850, when both compared against ERA5. Blue indicates better performance observed at this grid point, red indicates worse performance observed at this grid point.
C.4 Effects of Lateral Boundary Conditions
We run a scaled down toy model on different $W$ sizes of the Lateral Boundary Conditions. Larger $W$ yields better performance. See Figure 36
Figure 36: Root Mean Squared Error difference between different LBC sizes.
C.5 Spectra
We follow WeatherBench2[34] and adopt its zonal energy spectrum calculation in a limited area setting. See power spectrum analysis of 8 headline variables in Figure 37. The DLWP-LAMs have markedly better finer detail retention than its host model FourCastNet(SFNO), with MetMamba taking first place among majority of evaluated variables.
(a) 2t
(b) 10v
(c) msl
(d) z@500
(e) t@850
(f) q@700
(g) u@850
(h) v@850
Figure 37: Power spectrum of different DLWP models at forecast lead time 120 hour